What energy in kcal is needed to change 15.0g of ice at -10°c to 15.0g of water at 0°c?


1
Expert's answer
2021-04-16T05:00:45-0400

The specific heat capacity of ice is 2.05 J/goC

The heat of fusion of water is 333.55 J/g

Therefore, the total heat equals:

"Q_{total}=c_{ice}m\Delta{T}+c_fm=2.05J/g\cdot\degree{C}\times15.0g\times(0\degree{C}-(-10\degree{C}))+333.55J/g\times15.0g=5310.75J"


"5310.75J\times\frac{1kJ}{1000J}\times\frac{1kcal}{4.1868kJ}=1.27kcal"


Answer: 1.27 kcal